OpenStudy (anonymous):

The total cost for a bucket of popcorn and 4 movie tickets is $56. The total cost for the same size bucket of popcorn and 6 movie tickets is $80. Which equation represents the relationship between y, the total cost of the popcorn and movie tickets, and x, the number of movie tickets that are purchased?

OpenStudy (amaris6919):

what lesson is this.

OpenStudy (anonymous):

linear equations

OpenStudy (texaschic101):

y = popcorn and x = ticket y + 4x = 56 y + 6x = 80 these are the equations you will use...do you prefer substitution or elimination ?

OpenStudy (texaschic101):

oh...read the q wrong....do you have answer choices ?

OpenStudy (amaris6919):

yes the answer choices will be great can you show them

OpenStudy (anonymous):

y=12x+8 y=12x+24 y=4x+6 y=14x+28

OpenStudy (texaschic101):

(4,56)...x1 = 4 and y1 = 56 (6,80)...x2 = 6 and y2 = 80 slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (80 - 56) / (6 - 4) slope(m) = 24/2 slope(m) = 12 y = mx + b slope(m) = 12 (4,56)...x = 4 and y = 56 56 = 12(4) + b 56 = 48 + b 56 - 48 = b 8 = b y = 12x + 8 <==answer

OpenStudy (amaris6919):

y= 12x + 8 is the answer